OpenStudy (anonymous):

the length of a rectangular garden is 3m greater than the width. The area is 70m squared. find the dimensions of the garden.

OpenStudy (anonymous):

l=3+w a=l*w=70 70=l*w 70=(3+w)w 70=w^2+3w w^2+3w-70=0 (w+10)(w-7) w=-10 or 7 7 is the only valid width so w=7 l=w+3=7+3=10 so w=7 l=10